# Prediction Instability in Machine Learning Ensembles

### v1

@@ Acknowledgements
where $s(a)\backslash s(b)$ denotes any of $s(\cdot,a)>s(\cdot,b)$, $s(\cdot,a)<s(\cdot,b)$ or $s(\cdot,a)=s(\cdot,b)$. Note that for any $s^{*}\in S^{*}$ we have $s^{*}_{j}(\cdot,y)>s_{j}^{*}(\cdot,y^{\prime})$ if and only if $s_{j}^{j^{*}}(\cdot,y)>s_{j}^{j^{*}}(\cdot,y^{\prime})$ for all $j$ and so by Lemma 3.3 we have $y^{\prime}\in C(\{y,y^{\prime}\},s^{*})$. Moreover, for any $s^{*}\in S^{*}$ we have that $s^{*}_{j}(\cdot,y)>s_{j}^{*}(\cdot,y^{\prime\prime})$ if and only if the same holds in 9 for all $j$ and so Lemma 3.3 also implies that $y=C(\{y,y^{\prime\prime}\},s^{*})$. Thus by Lemma 5.1 we have $y^{\prime}=C(\{y^{\prime},y^{\prime\prime}\},s^{*})$. But now note that for any $s$ such that $s_{j^{*}}(\cdot,y^{\prime})>s_{j^{*}}(\cdot,y^{\prime\prime})$ there exists $s^{*}\in S^{*}$ such that $s_{j}(\cdot,y^{\prime})>s_{j}(\cdot,y^{\prime\prime})$ if and only if $s^{*}_{j}(\cdot,y^{\prime})>s^{*}_{j}(\cdot,y^{\prime\prime})$. By Lemma 3.3 we have $y^{\prime}=C(\{y^{\prime},y^{\prime\prime}\},s)$. We conclude that $j^{*}$ is decisive for $y^{\prime}$ over $y^{\prime\prime}$.
@@
Step 3: for all $\mathbf{y^{\prime\prime}\neq y,y^{\prime}}$ we have $\mathbf{j^{*}}$ decisive for $\mathbf{y}$ over $\mathbf{y^{\prime\prime}}$. Consider the set of ensemble scores $S^{*}\subseteq S(Y)$ such that for all $s^{*}\in S^{*}$:
$\begin{array}[]{lll}s^{*}_{1}(\cdot,y)\backslash s^{*}_{1}(\cdot,y^{\prime\prime})&>s^{*}_{1}(\cdot,y^{\prime})&>\ldots\\ \vdots&\vdots&\ddots\\ s^{*}_{j^{*}}(\cdot,y)&>s^{*}_{j^{*}}(\cdot,y^{\prime})&>s^{*}_{j^{*}}(\cdot,y^{\prime\prime})>\ldots\\ s^{*}_{j^{*}+1}(\cdot,y)&>s^{*}_{j^{*}+1}(\cdot,y^{\prime})&>\ldots\\ \hskip 19.91692pt\backslash s^{*}_{j^{*}+1}(\cdot,y^{\prime\prime})&&\\ \vdots&\vdots&\ddots\\ s^{*}_{m}(\cdot,y)\backslash s^{*}_{m}(\cdot,y^{\prime\prime})&>s^{*}_{m}(\cdot,y^{\prime})&>\ldots\\ \end{array}$ (26)
@@
By step (2) we have that $j^{*}$ is decisive for $y^{\prime}$ over $y^{\prime\prime}$ and so $y^{\prime}=C(\{y^{\prime},y^{\prime\prime}\},s^{*})$. By ensemble unanimity we have $y=C(\{y,y^{\prime}\},s^{*})$. By Lemma 5.1 we have $y=C(\{y,y^{\prime\prime}\},s^{*})$. But now note that for any $s$ such that $s_{j^{*}}(\cdot,y)>s_{j^{*}}(\cdot,y^{\prime\prime})$ there exists $s^{*}\in S^{*}$ such that $s_{j}(\cdot,y)>s_{j}(\cdot,y^{\prime\prime})$ if and only if $s^{*}_{j}(\cdot,y)>s_{j}^{*}(\cdot,y^{\prime\prime})$ for all $j$ and so by Lemma 3.3 we have that $y=C(\{y,y^{\prime\prime}\},s)$. We conclude that $j^{*}$ is decisive for $y$ over $y^{\prime\prime}$.
@@
$\begin{array}[]{lllll}s_{1}(\cdot,y^{\prime})&>s_{1}(\cdot,y^{\prime\prime})&>s_{1}(\cdot,y)&\ldots\\ \vdots&\vdots&\vdots&\ddots\\ s_{j^{*}}(\cdot,y^{\prime\prime})&>s_{j^{*}}(\cdot,y)&>s_{j^{*}}(\cdot,y^{\prime})&\ldots\\ \vdots&\vdots&\vdots&\ddots\\ s_{m}(\cdot,y^{\prime\prime})&>s_{m}(\cdot,y)&>s_{m}(\cdot,y^{\prime})&\ldots\\ \end{array}$ (32)
@@
Note that $s_{j}(\cdot,y)>s_{j}(\cdot,y^{\prime})$ iff $s^{j^{*}-1}_{j}(\cdot,y)>s^{j^{*}-1}_{j}(\cdot,y^{\prime})$ for all $j$ and so by Lemma 3.3 we have $y\in C(\{y,y^{\prime}\},s)$. By ensemble unanimity $y^{\prime\prime}=C(\{y,y^{\prime\prime}\},s)$ and so by Lemma 5.1 we have $y^{\prime\prime}=C(\{y^{\prime},y^{\prime\prime}\},s)$. Consider the set of ensemble scores $S^{*}\subseteq S(Y)$ such that for all $s^{*}\in S^{*}$:
$\begin{array}[]{lll}s^{*}_{1}(\cdot,y^{\prime})&>s^{*}_{1}(\cdot,y)&>\ldots\\ &\hskip 19.91692pt\backslash s^{*}_{1}(\cdot,y^{\prime\prime})&\\ \vdots&\vdots&\ddots\\ s^{*}_{j^{*}}(\cdot,y^{\prime\prime})&>s^{*}_{j^{*}}(\cdot,y^{\prime})&>s^{*}_{j^{*}}(\cdot,y)\geq\ldots\\ s^{*}_{j^{*}+1}(\cdot,y)&>s^{*}_{j^{*}+1}(\cdot,y^{\prime})&>\ldots\\ \hskip 19.91692pt\backslash s^{*}_{j^{*}+1}(\cdot,y^{\prime\prime})&&\\ \vdots&\vdots&\ddots\\ s^{*}_{m}(\cdot,y)&>s^{*}_{m}(\cdot,y^{\prime})&>\ldots\\ \hskip 19.91692pt\backslash s^{*}_{m}(\cdot,y^{\prime\prime})&&\end{array}$ (42)
@@
Note that for any $s^{*}\in S^{*}$ we have $s^{*}_{j}(\cdot,y)>s^{*}_{j}(\cdot,y^{\prime})$ iff $s^{j^{*}}_{j}(\cdot,y)>s^{j^{*}}_{j}(\cdot,y^{\prime})$ and so by Lemma 3.3 we have $y^{\prime}\in C(\{y,y^{\prime}\},s^{*})$. Also, $s^{*}_{j}(\cdot,y^{\prime})>s^{*}_{j}(\cdot,y^{\prime\prime})$ iff the same holds in 32 and so by Lemma 3.3 we have $y^{\prime\prime}=C(\{y^{\prime},y^{\prime\prime}\},s^{*})$. By Lemma 5.1 it follows that $y^{\prime\prime}=C(\{y,y^{\prime\prime}\},s^{*})$. But now note that for any $s$ such that $s_{j^{*}}(\cdot,y^{\prime\prime})>s_{j^{*}}(\cdot,y)$ there exists $s^{*}\in S^{*}$ such that $s_{j}(\cdot,y^{\prime\prime})>s_{j}(\cdot,y)$ iff $s_{j}^{*}(\cdot,y^{\prime\prime})>s_{j}^{*}(\cdot,y)$ for all $j$ and so by Lemma 3.3 we have $y^{\prime\prime}=C(\{y,y^{\prime\prime}\},s)$. We conclude that $j^{*}$ is decisive for $y^{\prime\prime}$ over $y$.
@@
Step 5: for all $\mathbf{y^{\prime\prime}\neq y,y^{\prime}}$ we have $\mathbf{j^{*}}$ decisive for $\mathbf{y^{\prime\prime}}$ over $\mathbf{y^{\prime}}$. Consider the set of ensemble scores $S^{*}\subseteq S(Y)$ such that for all $s^{*}\in S^{*}$:
$\begin{array}[]{lll}s^{*}_{1}(\cdot,y)&>s^{*}_{1}(\cdot,y^{\prime})&>\ldots\\ &\hskip 19.91692pt\backslash s^{*}_{1}(\cdot,y^{\prime\prime})&\\ \vdots&\vdots&\ddots\\ s^{*}_{j^{*}}(\cdot,y^{\prime\prime})&>s^{*}_{j^{*}}(\cdot,y)&>s^{*}_{j^{*}}(\cdot,y^{\prime})>\ldots\\ s^{*}_{j^{*}+1}(\cdot,y)&>s^{*}_{j^{*}+1}(\cdot,y^{\prime})&>\ldots\\ &\hskip 19.91692pt\backslash s^{*}_{j^{*}+1}(\cdot,y^{\prime\prime})&\\ \vdots&\vdots&\ddots\\ s^{*}_{m}(\cdot,y)&>s^{*}_{m}(\cdot,y^{\prime})&>\ldots\\ &\hskip 19.91692pt\backslash s^{*}_{m}(\cdot,y^{\prime\prime})&\end{array}$ (52)
@@
By step (4) we have that $j^{*}$ is decisive for $y^{\prime\prime}$ over $y$, and so $y^{\prime\prime}=C(\{y,y^{\prime\prime}\},s)$. By ensemble unanimity $y=C(\{y,y^{\prime}\},s^{*})$. By Lemma 5.1 we have $y^{\prime\prime}=C(\{y^{\prime},y^{\prime\prime}\},s^{*})$. But now note that for any $s$ such that $s_{j^{*}}(\cdot,y^{\prime\prime})>s_{j^{*}}(\cdot,y^{\prime})$ there exists $s^{*}\in S^{*}$ such that $s_{j}(\cdot,y^{\prime\prime})>s_{j}(\cdot,y^{\prime})$ iff $s^{*}_{j}(\cdot,y^{\prime\prime})>s^{*}_{j}(\cdot,y^{\prime})$ for all $j$ and so by Lemma 3.3 we have that $y^{\prime\prime}=C(\{y^{\prime},y^{\prime\prime}\},s)$. We conclude that $j^{*}$ is decisive for $y^{\prime\prime}$ over $y^{\prime}$.
@@
$\begin{array}[]{lll}s^{*}_{1}(\cdot,y)&>&\ldots\\ \hskip 19.91692pt\backslash s^{*}_{1}(\cdot,y^{\prime})&&\\ \vdots&\vdots&\ddots\\ s^{*}_{j^{*}}(\cdot,y)&>s^{*}_{j^{*}}(\cdot,y^{\prime\prime})&>s^{*}_{j^{*}}(\cdot,y^{\prime})>\ldots\\ s^{*}_{j^{*}+1}(\cdot,y)&>&\ldots\\ \hskip 19.91692pt\backslash s^{*}_{j^{*}+1}(\cdot,y^{\prime})&&\\ \vdots&\vdots&\ddots\\ s^{*}_{m}(\cdot,y)&>&\ldots\\ \hskip 19.91692pt\backslash s^{*}_{m}(\cdot,y^{\prime})&&\end{array}$ (62)
@@
By step (5) we have that $j^{*}$ is decisive for $y^{\prime\prime}$ over $y^{\prime}$ and so $y^{\prime\prime}=C(\{y^{\prime},y^{\prime\prime}\},s^{*})$. By step (3) we have that $j^{*}$ is decisive for $y$ over $y^{\prime\prime}$ and so $y=C(\{y,y^{\prime\prime}\},s^{*})$. By Lemma 5.1 we have $y=C(\{y,y^{\prime}\},s^{*})$. But now note that for any $s$ such that $s_{j^{*}}(\cdot,y)>s_{j^{*}}(\cdot,y^{\prime})$ there exists $s^{\prime}\in S^{*}$ such that $s_{j}(\cdot,y)>s_{j}(\cdot,y^{\prime})$ iff $s^{*}_{j}(\cdot,y)>s^{*}_{j}(\cdot,y^{\prime})$ and so by Lemma 3.3 we have $y=C(\{y,y^{\prime}\},s)$. We conclude that $j^{*}$ is decisive for $y$ over $y^{\prime}$. The argument for $j^{*}$ decisive for $y^{\prime}$ over $y$ is the same with the two reversed, citing step (4) and step (2).
@@
Step 7: $\mathbf{j^{*}}$ is decisive for any $a\in Y$ over any $b\in Y$. Take any $a\neq y,y^{\prime}$ and any $b\neq y,y^{\prime}$ and consider the set of scores $S^{*}\subseteq S(Y)$ such that for all $s^{*}\in S^{*}$:
$\begin{array}[]{lll}s^{*}_{1}(\cdot,a)&>s^{*}_{1}(\cdot,y)&\ldots\\ \hskip 19.91692pt\backslash s^{*}_{1}(\cdot,b)&\hskip 19.91692pt\backslash s^{*}_{1}(\cdot,y^{\prime})&\\ \vdots&\vdots&\ddots\\ s^{*}_{j^{*}}(\cdot,a)&>s^{*}_{j^{*}}(\cdot,y)&>s^{*}_{j^{*}}(\cdot,y^{\prime})>s^{*}_{j^{*}}(\cdot,b)\ldots\\ \vdots&\vdots&\ddots\\ s^{*}_{m}(\cdot,a)&>s^{*}_{m}(\cdot,y)&\ldots\\ \hskip 19.91692pt\backslash s^{*}_{m}(\cdot,b)&\hskip 19.91692pt\backslash s^{*}_{m}(\cdot,y^{\prime})&\end{array}$ (70)
@@
By step (4) we have that $j^{*}$ is decisive for $a$ over $y$, by step (6) we have that $j^{*}$ decisive for $y$ over $y^{\prime}$, and by step (2) we have that $j$ is decisive for $y^{\prime}$ over $b$. Thus by Lemma 5.1 we have $a=C(\{a,b\},s^{*})$. But now note that for any $s$ such that $s_{j^{*}}(\cdot,a)>s_{j^{*}}(\cdot,b)$ there exists $s^{*}\in S^{*}$ such that $s_{j}(\cdot,a)>s_{j}(\cdot,b)$ iff $s^{*}_{j}(\cdot,a)>s^{*}_{j}(\cdot,b)$ for all $j$, and so by Lemma 3.3 we have $a=C(\{a,b\},s)$. Thus $j^{*}$ decisive for $a$ over $b$ and since $a$ and $b$ were arbitrary, there exists $j$, namely $j^{*}$, such that for all $y,y^{\prime}\in Y$ if $s_{j^{*}}(\cdot,y)>s_{j^{*}}(\cdot,y^{\prime})$ then $y=C(\{y,y^{\prime}\},s)$. Thus definition 2.2 is violated, a contradiction. ∎
Lemma 5.2.
If $C(\cdot,s)$ is monotonic then it respects model choice reversal.
@@
Take any $C(\cdot)$ and suppose that it is monotonic. Take any $y,y^{\prime}\in Y$ and any $s,s^{\prime}$ such that $y\in C(\{y,y^{\prime}\},s)$, $y^{\prime}\notin C(\{y,y^{\prime}\},s)$, and $y^{\prime}\in C(\{y,y^{\prime}\},s^{\prime})$. Suppose by contradiction that there does not exist $j$ such that $s_{j}(\cdot,y)>s_{j}(\cdot,y^{\prime})$ and $s^{\prime}_{j}(\cdot,y)<s_{j}^{\prime}(\cdot,y^{\prime})$. That is, for all $j$ if $s_{j}(\cdot,y)>s_{j}(\cdot,y^{\prime})$ then $s_{j}^{\prime}(\cdot,y)\geq s_{j}^{\prime}(\cdot,y^{\prime})$ and if $s_{j}^{\prime}(\cdot,y)<s_{j}^{\prime}(\cdot,y^{\prime})$ then $s_{j}(\cdot,y)\geq s_{j}(\cdot,y^{\prime})$. But then $s_{j}^{\prime}(\cdot,y^{\prime})\geq s_{j}^{\prime}(\cdot,y)$ implies that $s_{j}(\cdot,y^{\prime})\geq s_{j}(\cdot,y)$ for all $j$. By monotonicity $y^{\prime}\in C(\{y,y^{\prime}\},s^{\prime})$ implies that $y^{\prime}\in C(\{y,y^{\prime}\},s)$, a contradiction. ∎
Proof of Corollary 4.4.
@@
Suppose that $|Y|\geq 3$, that $C(Y^{*},s)\neq\emptyset$ for all nonempty $Y^{*}\subseteq Y$, and that $C(\cdot,s)$ is insertion/deletion consistent. Then by the ECD theorem there exists $S^{*}\subseteq S(Y)$ such that at least one of definitions 2.2, 2.3, or 2.4 are violated on $S^{*}$. By $s(\cdot)$ surjective, for all $s^{*}\in S^{*}$ there exists $x\in X$ such that $s(x)=s^{*}$. Thus, define:
$X^{*}\equiv\bigcup_{s^{*}\in S^{*}}\left\{x\in X|s(x)=s^{*}\right\}.$

### v5

@@ Acknowledgment
where $z_{j,d(a)}\backslash z_{j,d(b)}$ denotes any of $z_{j,d(a)}>z_{j,d(b)}$, $z_{j,d(a)}<z_{j,d(b)}$ or $z_{j,d(a)}=z_{j,d(b)}$. Note that for any $z^{*}\in S^{*}$ we have $z^{*}_{j,d(y)}>z^{*}_{j,d(y^{\prime})}$ if and only if $z^{j^{*}}_{j,d(y)}>z^{j^{*}}_{j,d(y^{\prime})}$ for all $j$ and so by Lemma III.3 we have $y^{\prime}\in C(\{y,y^{\prime}\},z^{*})$. Moreover, for any $z^{*}\in S^{*}$ we have that $z^{*}_{j,d(y)}>z^{*}_{j,d(y^{\prime\prime})}$ if and only if the same holds in 4 for all $j$ and so Lemma III.3 also implies that $\{y\}=C(\{y,y^{\prime\prime}\},z^{*})$. Thus by Lemma VI.1 we have $\{y^{\prime}\}=C(\{y^{\prime},y^{\prime\prime}\},z^{*})$. But now note that for any $z$ such that $z_{j^{*},d(y^{\prime})}>z_{j^{*},d(y^{\prime\prime})}$ there exists $z^{*}\in S^{*}$ such that $z_{j,d(y^{\prime})}>z_{j,d(y^{\prime\prime})}$ if and only if $z^{*}_{j,d(y^{\prime})}>z^{*}_{j,d(y^{\prime\prime})}$. By Lemma III.3 we have $\{y^{\prime}\}=C(\{y^{\prime},y^{\prime\prime}\},z)$. We conclude that $j^{*}$ is decisive for $y^{\prime}$ over $y^{\prime\prime}$.
@@
Step 3: for all $\mathbf{y^{\prime\prime}\neq y,y^{\prime}}$ we have $\mathbf{j^{*}}$ decisive for $\mathbf{y}$ over $\mathbf{y^{\prime\prime}}$. Consider the set of ensemble scores $S^{*}\subseteq S(Y)$ such that for all $z^{*}\in S^{*}$:
$z^{*}_{1,d(y)}\backslash z^{*}_{1,d(y^{\prime\prime})}$ $>z^{*}_{1,d(y^{\prime})}>\ldots$
$\vdots$
$z^{*}_{j^{*},d(y)}$ $>z^{*}_{j^{*},d(y^{\prime})}>z^{*}_{j^{*},d(y^{\prime\prime})}>\ldots$
$z^{*}_{j^{*}+1,d(y)}\backslash z^{*}_{j^{*}+1,d(y^{\prime\prime})}$ $>z^{*}_{j^{*}+1,d(y^{\prime})}>\ldots$
$\vdots$
$z^{*}_{m,d(y)}\backslash z^{*}_{m,d(y^{\prime\prime})}$ $>z^{*}_{m,d(y^{\prime})}>\ldots$ (6)
@@
By step (2) we have that $j^{*}$ is decisive for $y^{\prime}$ over $y^{\prime\prime}$ and so $\{y^{\prime}\}=C(\{y^{\prime},y^{\prime\prime}\},z^{*})$. By ensemble unanimity we have $\{y\}=C(\{y,y^{\prime}\},z^{*})$. By Lemma VI.1 we have $\{y\}=C(\{y,y^{\prime\prime}\},z^{*})$. But now note that for any $z$ such that $z_{j^{*},d(y)}>z_{j^{*},d(y^{\prime\prime})}$ there exists $z^{*}\in S^{*}$ such that $z_{j,d(y)}>z_{j,d(y^{\prime\prime})}$ if and only if $z^{*}_{j,d(y)}>z^{*}_{j,d(y^{\prime\prime})}$ for all $j$ and so by Lemma III.3 we have that $\{y\}=C(\{y,y^{\prime\prime}\},z)$. We conclude that $j^{*}$ is decisive for $y$ over $y^{\prime\prime}$.
@@
$z_{1,d(y^{\prime})}$ $>z_{1,d(y^{\prime\prime})}>z_{1,d(y)}\ldots$
$\vdots$
$z_{j^{*},d(y^{\prime\prime})}$ $>z_{j^{*},d(y)}>z_{j^{*},d(y^{\prime})}\ldots$
$\vdots$
$z_{m,d(y^{\prime\prime})}$ $>z_{m,d(y)}>z_{m,d(y^{\prime})}\ldots$ (7)
@@
Note that $z_{j,d(y)}>z_{j,d(y^{\prime})}$ iff $z^{j^{*}-1}_{j,d(y)}>z^{j^{*}-1}_{j,d(y)}$ for all $j$ and so by Lemma III.3 we have $y\in C(\{y,y^{\prime}\},z)$. By ensemble unanimity $\{y^{\prime\prime}\}=C(\{y,y^{\prime\prime}\},z)$ and so by Lemma VI.1 we have $\{y^{\prime\prime}\}=C(\{y^{\prime},y^{\prime\prime}\},z)$. Consider the set of ensemble scores $S^{*}\subseteq S(Y)$ such that for all $z^{*}\in S^{*}$:
$z^{*}_{1,d(y^{\prime})}$ $>z^{*}_{1,d(y)}\backslash z^{*}_{1,d(y^{\prime\prime})}>\ldots$
$\vdots$
$z^{*}_{j^{*},d(y^{\prime\prime})}$ $>z^{*}_{j^{*},d(y^{\prime})}>z^{*}_{j^{*},d(y)}\geq\ldots$
$z^{*}_{j^{*}+1,d(y)}\backslash z^{*}_{j^{*}+1,d(y^{\prime\prime})}$ $>z^{*}_{j^{*}+1,d(y^{\prime})}>\ldots$
$\vdots$
$z^{*}_{m,d(y)}\backslash z^{*}_{m,d(y^{\prime\prime})}$ $>z^{*}_{m,d(y^{\prime})}>\ldots$
@@
Note that for any $z^{*}\in S^{*}$ we have $z^{*}_{j,d(y)}>z^{*}_{j,d(y^{\prime})}$ iff $z^{j^{*}}_{j,d(y)}>z^{j^{*}}_{j,d(y^{\prime})}$ and so by Lemma III.3 we have $y^{\prime}\in C(\{y,y^{\prime}\},z^{*})$. Also, $z^{*}_{j,d(y^{\prime})}>z^{*}_{j,d(y^{\prime\prime})}$ iff the same holds in 7 and so by Lemma III.3 we have $\{y^{\prime\prime}\}=C(\{y^{\prime},y^{\prime\prime}\},z^{*})$. By Lemma VI.1 it follows that $\{y^{\prime\prime}\}=C(\{y,y^{\prime\prime}\},z^{*})$. But now note that for any $z$ such that $z_{j^{*},d(y^{\prime\prime})}>z_{j^{*},d(y)}$ there exists $z^{*}\in S^{*}$ such that $z_{j,d(y^{\prime\prime})}>z_{j,d(y)}$ iff $z^{*}_{j,d(y^{\prime\prime})}>z^{*}_{j,d(y)}$ for all $j$ and so by Lemma III.3 we have $\{y^{\prime\prime}\}=C(\{y,y^{\prime\prime}\},z)$. We conclude that $j^{*}$ is decisive for $y^{\prime\prime}$ over $y$.
@@
Step 5: for all $\mathbf{y^{\prime\prime}\neq y,y^{\prime}}$ we have $\mathbf{j^{*}}$ decisive for $\mathbf{y^{\prime\prime}}$ over $\mathbf{y^{\prime}}$. Consider the set of ensemble scores $S^{*}\subseteq S(Y)$ such that for all $z^{*}\in S^{*}$:
$z_{1,d(y)}^{*}$ $>z_{1,d(y^{\prime})}^{*}\backslash z_{1,d(y^{\prime\prime})}^{*}>\ldots$
$\vdots$
$z_{j^{*},d(y^{\prime\prime})}^{*}$ $>z_{j^{*},d(y)}^{*}>z_{j^{*},d(y^{\prime})}^{*}>\ldots$
$z_{j^{*}+1,d(y)}^{*}$ $>z_{j^{*}+1,d(y^{\prime})}^{*}\backslash z_{j^{*}+1,d(y^{\prime\prime})}^{*}>\ldots$
$\vdots$
$z_{m,d(y)}^{*}$ $>z_{m,d(y^{\prime})}^{*}\backslash z_{m,d(y^{\prime\prime})}^{*}>\ldots$ (8)
@@
By step (4) we have that $j^{*}$ is decisive for $y^{\prime\prime}$ over $y$, and so $\{y^{\prime\prime}\}=C(\{y,y^{\prime\prime}\},z)$. By ensemble unanimity $\{y\}=C(\{y,y^{\prime}\},z^{*})$. By Lemma VI.1 we have $\{y^{\prime\prime}\}=C(\{y^{\prime},y^{\prime\prime}\},z^{*})$. But now note that for any $z$ such that $z_{j^{*},d(y^{\prime\prime})}>z_{j^{*},d(y^{\prime})}$ there exists $z^{*}\in S^{*}$ such that $z_{j,d(y^{\prime\prime})}>z_{j,d(y^{\prime})}$ iff $z^{*}_{j,d(y^{\prime\prime})}>z^{*}_{j,d(y^{\prime})}$ for all $j$ and so by Lemma III.3 we have that $\{y^{\prime\prime}\}=C(\{y^{\prime},y^{\prime\prime}\},z)$. We conclude that $j^{*}$ is decisive for $y^{\prime\prime}$ over $y^{\prime}$.
@@
$z^{*}_{1,d(y)}\backslash z^{*}_{1,d(y^{\prime})}$ >$\ldots$> …
$\vdots$
$z^{*}_{j^{*},d(y)}$ $>z^{*}_{j^{*},d(y^{\prime\prime})}>z^{*}_{j^{*},d(y^{\prime})}>\ldots$
$z^{*}_{j^{*}+1,d(y)}\backslash z^{*}_{j^{*}+1,d(y^{\prime})}$ >$\ldots$> …
$\vdots$
$z^{*}_{m,d(y)}\backslash z^{*}_{m,d(y^{\prime})}$ >$\ldots$> … (9)
@@
By step (5) we have that $j^{*}$ is decisive for $y^{\prime\prime}$ over $y^{\prime}$ and so $\{y^{\prime\prime}\}=C(\{y^{\prime},y^{\prime\prime}\},z^{*})$. By step (3) we have that $j^{*}$ is decisive for $y$ over $y^{\prime\prime}$ and so $\{y\}=C(\{y,y^{\prime\prime}\},z^{*})$. By Lemma VI.1 we have $\{y\}=C(\{y,y^{\prime}\},z^{*})$. But now note that for any $z$ such that $z_{j^{*},d(y)}>z_{j^{*},d(y^{\prime})}$ there exists $z^{*}\in S^{*}$ such that $z_{j,d(y)}>z_{j,d(y^{\prime})}$ iff $z^{*}_{j,d(y)}>z^{*}_{j,d(y^{\prime})}$ and so by Lemma III.3 we have $\{y\}=C(\{y,y^{\prime}\},z)$. We conclude that $j^{*}$ is decisive for $y$ over $y^{\prime}$. The argument for $j^{*}$ decisive for $y^{\prime}$ over $y$ is the same with the two reversed, citing step (4) and step (2).
@@
Step 7: $\mathbf{j^{*}}$ is decisive for any $a\in Y$ over any $b\in Y$. Take any $a\neq y,y^{\prime}$ and any $b\neq y,y^{\prime}$ and consider the set of scores $S^{*}\subseteq S(Y)$ such that for all $z^{*}\in S^{*}$:
$z^{*}_{1,d(a)}\backslash z^{*}_{1,d(b)}$ $>z^{*}_{1,d(y)}\backslash z^{*}_{1,d(y^{\prime})}\ldots$
$\vdots$
$z^{*}_{j^{*},d(a)}$ $>z^{*}_{j^{*},d(y)}>z^{*}_{j^{*},d(y^{\prime})}>z^{*}_{j^{*},d(b)}\ldots$
$\vdots$
$z^{*}_{m,d(a)}\backslash z^{*}_{m,d(b)}$ $>z^{*}_{m,d(y)}\backslash z^{*}_{m,d(y^{\prime})}\ldots$ (10)
@@
By step (4) we have that $j^{*}$ is decisive for $a$ over $y$, by step (6) we have that $j^{*}$ decisive for $y$ over $y^{\prime}$, and by step (2) we have that $j$ is decisive for $y^{\prime}$ over $b$. Thus by Lemma VI.1 we have $\{a\}=C(\{a,b\},z^{*})$. But now note that for any $z$ such that $z_{j^{*},d(a)}>z_{j^{*},d(b)}$ there exists $z^{*}\in S^{*}$ such that $z_{j,d(a)}>z_{j,d(b)}$ iff $z^{*}_{j,d(a)}>z^{*}_{j,d(b)}$ for all $j$, and so by Lemma III.3 we have $\{a\}=C(\{a,b\},z)$. Thus $j^{*}$ decisive for $a$ over $b$ and since $a$ and $b$ were arbitrary, there exists $j$, namely $j^{*}$, such that for all $y,y^{\prime}\in Y$ if $z_{j^{*},d(y)}>z_{j^{*},d(y^{\prime})}$ then $\{y\}=C(\{y,y^{\prime}\},z)$. Thus definition II.5 is violated, a contradiction. This establishes (1).
By the forgoing, we have established that any $C(\cdot)$ that satisfies definition II.5 at least one of definitions II.4, II.6, or II.7. Thus, if $C(\cdot)$ that satisfies definition II.5 there exists $S^{*}\subseteq S(Y)$ such that at least one of definitions II.4, II.6, or II.7 are violated on $S^{*}$. By $s(\cdot)$ surjective, for all $s^{*}\in S^{*}$ there exists $x\in X$ such that $s(x)=s^{*}$. Thus, define:
$X^{*}\equiv\bigcup_{s^{*}\in S^{*}}\left\{x\in X|s(x)=s^{*}\right\}.$
This establishes (2).
Given a strict ordering $P\in\mathbf{R}$ define:
$\overline{W}(Y,P)=\left\{z\in\mathbb{R}^{|Y|}\left|\mbox{for all }y,y^{\prime}\in Y\atop{y\hskip 2.84526ptP\hskip 2.84526pty^{\prime}\mbox{ iff}z_{d(y)}>z_{d(y^{\prime})}}\right.\right\}.$
Lemma VI.2.
Consider a restriction on $S(Y)$ such that for any $z\in S(Y)$ there exists $P\in\mathbf{R}$ such that $z_{j}\in\overline{W}(Y,P)$ for all $j$. Then if $C(\cdot)$ satisfies ensemble unanimity it respects model choice reversal.
@@
Assume that the hypotheses of the lemma hold and consider any $C(\cdot)$, any $z,z^{\prime}\in S(Y)$, and any $y,y^{\prime}\in Y$ such that $y=C(\{y,y^{\prime}\},z)$ and $y^{\prime}\in C(\{y,y^{\prime}\},z^{\prime})$. By assumption, there exists $P,P^{\prime}\in\mathbf{R}$ such that for all $j$ we have $z_{j}\in\overline{W}(Y,P)$ and $z^{\prime}\in\overline{W}(Y,P^{\prime})$. Thus, it must be that for all $j$ we have $z_{j,d(y)}>z_{j,d(y^{\prime})}$, else by unanimity $y^{\prime}\in C(\{y,y^{\prime}\},z)$, a contradiction. Suppose that there does not exist $j$ such that $z_{j,d(y)}>z_{j,d(y^{\prime})}$ and $z^{\prime}_{j,d(y)}<z^{\prime}_{j,d(y^{\prime})}$. Then $z^{\prime}\in\overline{W}(Y,P^{\prime})$ implies that $z^{\prime}_{j,d(y)}>z^{\prime}_{j,d(y^{\prime})}$ for all $j$. By ensemble unanimity $y=C(\{y,y^{\prime}\},z^{\prime})$, a contradiction. ∎
Proof of theorem V.2.
@@
Fix a finite set of label values $Y=\{y_{1},y_{2},\ldots\}$ and a true conditional distribution $p(Y=y|X=x)$. Generically, for any $x\in X$, there exists a strict ordering $P_{x}\in\mathbf{R}$ so that $\{p(y|x)\}_{y\in Y}\in\overline{W}(Y,P_{x})$. By $s_{j}(x|D_{n})$ fully consistent it follows that $s_{j}(x)\in\overline{W}(Y,P_{x})$ for all $j$ and all $x$ in the limit as $n\to\infty$ with probability one. Now, consider the soft voting ensemble choice aggregator:
$C(Y,s(x))=\operatornamewithlimits{argmax}_{y\in Y}\left\{\frac{1}{m}\sum_{j}s_{j}(x,y)\right\}.$
Clearly this non-degenerate and satisfies ensemble unanimity. It is also transitive and so by Lemma III.2 satisfies insertion/deletion consistency. Finally, by Lemma VI.2 it also satisfies model choice reversal for all $x\in X$ with probability one. ∎